# Trust Me, I Can Convince You: The Contextualized Argument Appraisal Framework

### v1

@@ 6.1 RQ1: How do discrete emotions and appraisals correlate with argument convincingness?
Table 3 displays the Pearson’s correlation values between all variables and convincingness. All correlations between emotions and convincingness are significant, except for fear. The strongest positive effects are with trust (.58), relief (.51), pride (.47), and joy (.45). In contrast, anger (-.22) and disgust (-.23) show negative correlations with convincingness. These findings align with Greschner and Klinger (2025).
@@
Appraisals.
We now focus on the analysis of emotions and appraisals in arguments, illustrated in Figure 4. For each emotion, the average appraisal value is displayed in the heatmap. The values are in line with other applications of appraisal theories to emotion analysis in natural language processing (Troiano et al., 2023, Figure 8), confirming the validity of the approach.
Unpleasantness, internal check and external check appraisals consistently show the highest ratings for anger (3.0, 2.8, 2.7) and disgust (3.2, 2.8, 2.9). Intuitively, arguments evoking negative emotions are being appraised as unpleasant, clashing with the receiver’s ideals, and violating social norms. In contrast, joy, pride, and relief show high values for pleasantness (3.5, 3.2, 3.0), familiarity (3.1, 3.3, 3.2), and positive consequentiality (2.5, 2.5, 2.5). This indicates that arguments being appraised as beneficial, familiar, and rewarding evoke positive emotions.
Beyond categorical emotions, appraisals also explain convincingness: we find significant correlations for all appraisal dimensions (see Table 3) except consequence manageability. Pleasantness shows a strong positive (.58), positive consequentiality a moderate (.40), and familiarity a moderate (.34) positive correlation with convincingness. In contrast, we find strong negative correlations with argument external check (-.50), unpleasantness (-.36), and external check (-.36).
Both discrete emotion and appraisal analyses demonstrate significant correlations between positive emotions and convincingness, and negative correlations between negative emotions and convincingness. These results align with Greschner and Klinger (2025), but contradict that fear increases convincingness (Dillard and Anderson, 2004).
@@
Table 4: Average values of how important the sender, the argument, and the receiver were for the emotional response of the receiver. The importance of the the 3 dimensions was rated on a 1–5 scale.
6.2 RQ2: Which contextual factors shape the emotional response to arguments?
@@
After analyzing which emotions are developed and how they are correlate with the argument’s convincingness, we now turn to understanding how these emotions are developed. We analyze the participant’s (argument receiver’s) textual explanations about the evoked emotions and the strongest influence on the emotion (sender, argument, receiver). We also analyze how strong (1–5) the influence (sender, argument, receiver) was for the development of the evoked emotion (see Table 4). The argument (3.6) is the primary driver of the emotional response, followed by the receiver themselves (3.1) and the sender (1.9). Receivers rarely attribute guilt and shame to being evoked by the sender.
@@
Pride is often attributed to the argument itself, which is also the most frequent emotion and has the highest correlation to convincingness. Similar patterns are found for anger, which has however has the strongest negative convincingness correlation. In summary, the argument itself is the primary driver of the emotional response.
7 Conclusion
@@
In this paper, we introduce the Contextualized Argument Appraisal Framework (CAAF), an approach to explain how arguments are cognitively appraised. In a synchronized process, individuals also develop an emotion and a perceive a particular convincingness. The analysis of the resulting corpus ContArgA, the data resulting from a role-playing annotation study based on CAAF, shows that appraisals are appropriate for explaining argument convincingness. We release the data, enabling research on arguments with focus on contextual (sender, receiver, argument) properties.
@@
Particularly, CAAF helps to disentangle the influences that shape a person’s reaction to arguments, because it captures the role of properties of the sender, receiver, and the argument itself. In downstream systems, these factors can contribute to argument mining and content moderation systems, for example to support studies on desinformation and argumentation strategies. The framework lays the groundwork for context-sensitive and emotionally aware computational systems that reflect how people engage with persuasive text.
@@
Regarding our study, each participant annotates two arguments from one topic. The argument selection from the topic is randomized, which can lead to biases due to the prior stance of the participant and the order of pro and con arguments. We acknowledge that future work can take a deep dive into the role of argument presentation according to stances, since we report the exact argument in displayed order within our available ContArgA corpus. Further, we allow participation up to 39 times, which ensures that one participant can only annotate two arguments per topic, however, there might be annotators that do in fact participate in all 39 studies (i.e, contributing 78 individual argument annotations out of the 4,000 arguments in ContArgA), therefore introducing annotator bias.
@@
The study is part of a larger project that has been approved, including a description of the presented study, by the ethics board of the University of [Anonymous]. Following recommendations with respect to ethical challenges in emotion analysis (Mohammad, 2022), we point out the following facts about our work: The purpose of the CAAF is to study the interplay of emotions in argument texts while taking contextual factors into account, focusing on the emotions evoked in receivers. We choose categorical emotion labels from emotion theories in accordance with previous work, and expand appraisal labels from dimensional appraisal theories to be applicable to arguments.
@@
Acknowledgments
This project has been conducted as part of the Emcona (The Interplay of Emotions and Convincingness in Arguments) project, which is funded by the German Research Foundation (DFG, project KL2869/12–1, project number 516512112). We thank Meike Bauer for her thorough work and the fruitful discussions.
@@ Acknowledgments
Appendix B Emotion Influence
@@
Appendix C Topics
@@
Appendix D Role-Playing Scenario
@@
Figure 6: Involvement question for the role-playing scenario in our annotation study.
Appendix E Study Design
The following section shows the screenshots of out study. After Figure 26 the study loops back to Figure 14 to display the second argument.

### v2

@@ 6.1.  RQ1: How do discrete emotions and appraisals correlate with argument convincingness?
Table 3 displays all correlation values of the variables of interest and convincingness with Bonferroni correction. All correlations between emotion and convincingness are significant, except for fear. The strongest positive effects are with trust (.57), relief (.51), pride (.46), and joy (.44). In contrast, anger (-.22) and disgust (-.23) show the strongest negative correlations with convincingness.
@@
Table 4: Average values of how important the sender, the argument, and the receiver are for the emotional response of the receiver. The importance of each dimension is rated on a 1–5 scale.
Appraisals.
We now focus on the analysis of emotions and appraisals in arguments, illustrated in Figure 4. For each emotion, the average appraisal value is displayed in the heatmap. To the best of our knowledge, we are the first to apply appraisal theory to arguments. However, the appraisal dimensions of suddenness, familiarity, pleasantness, unpleasantness, internal check, external check values are in line with other applications of appraisal theories to emotion analysis in natural language processing (Troiano et al., 2023, Figure 8), confirming the validity of the approach.
Turning to the analysis of appraisal dimensions in arguments, we find that unpleasantness, internal check and external check appraisals consistently show the highest ratings for anger (3.0, 2.8, 2.7) and disgust (3.2, 2.8, 2.9). Intuitively, arguments evoking negative emotions are being appraised as unpleasant, clashing with the receiver’s ideals, and violating social norms. In contrast, joy, pride, and relief show high values for pleasantness (3.5, 3.2, 3.0), familiarity (3.1, 3.3, 3.2), and positive consequentiality (2.5, 2.5, 2.5). This indicates that arguments being appraised as beneficial, familiar, and rewarding evoke positive emotions.
Beyond categorical emotions, appraisals are also related to convincingness, as displayed in Table 3. We find significant correlations for all appraisal dimensions except consequence manageability. Pleasantness shows a strong positive (.57), positive consequentiality (.39) and familiarity (.33) a moderate positive correlation with convincingness. In contrast, we find strong negative correlations with argument external check (-.50), unpleasantness (-.39), as well as external check (-.36).
Both discrete emotion and appraisal analyses demonstrate significant correlations between positive emotions and convincingness, and negative correlations between negative emotions and convincingness. The results with respect to discrete emotions align with Greschner and Klinger (2025) (showing negative correlations between negative emotions and convincingness using 300 German arguments), but contradict that fear increases convincingness (Dillard and Anderson, 2004).
6.2.  RQ2: Which contextual factors shape the emotional response to arguments?
@@
After analyzing which emotions are developed and how they are correlate with the argument’s convincingness, we now turn to understanding how these emotions are developed. We analyze the participant’s (argument receiver’s) textual explanations about the evoked emotions and the strongest influence on the emotion (sender, argument, receiver). We also analyze how strong (1–5) the influence (sender, argument, receiver) was for the development of the evoked emotion.
@@
Table 4 shows the results. The argument (3.6) is the primary driver of the emotional response, followed by the receiver themselves (3.1) and the sender (1.9). Receivers rarely attribute guilt and shame to being evoked by the sender. Pride is often attributed to the argument itself, which is the most frequent emotion and has the highest correlation with convincingness. Similar patterns are found for anger, which has the strongest negative correlation. Interestingly, the sender is the least important component when it comes to the evoked emotion of guilt. In summary, the argument itself is the primary driver of the emotional response.
7.  Conclusion
@@
In this paper, we introduce the Contextualized Argument Appraisal Framework (CAAF). It is a novel approach to explain how arguments are cognitively appraised. In a synchronized process, individuals develop an emotion and perceive a particular convincingness after argument exposure. The analysis of the resulting corpus ContArgA, the data resulting from a role-playing annotation study based on CAAF, shows that appraisals are appropriate for explaining the relationship between emotions and convincingness. Furthermore, our results validate the framework’s contextual approach: while the arguments themselves generally exert the strongest influence on emotional responses, emotion-specific patterns emerge regarding the relative contributions of sender, argument, and receiver factors. We release the data, enabling research on arguments with focus on such contextual properties.
@@
Particularly, CAAF helps to disentangle the influences that shape a person’s reaction to arguments, because it captures the role of properties of the sender, receiver, and the argument itself. These factors can contribute to argument mining and content moderation systems, for example to support studies on desinformation and argumentation strategies. The framework lays the groundwork for context-sensitive and emotionally aware computational systems that reflect how people engage with persuasive text.
@@
Regarding our study, each participant annotates two arguments from one topic. The argument selection from the topic is randomized, which can lead to biases due to the prior stance of the participant and the order of pro and con arguments. We acknowledge that future work can take a deep dive into the role of argument presentation according to stances, since we report the exact argument in displayed order within our available ContArgA corpus. Further, we allow participation up to 39 times, which ensures that one participant can only annotate two arguments per topic, however, there might be annotators that do in fact participate in all 39 studies (i.e, contributing 78 individual argument annotations out of the 4,000 arguments in ContArgA), therefore, introducing annotator bias.
To address the emerging challenge of automated bot participation in crowdsourcing studies, we manually verfy all free-text response to ensure authenticity. Participants whose completion times fell significantly below the study average were systematically excluded from analysis. Recognizing the inherent quality control challenges in crowdsourced research, we incorporated seven attention and saliency checks throughout the study. Only data from participants who successfully passed all validation measures were retained for analysis. These quality assurance protocols, while necessary to mitigate common crowdsourcing limitations, enabled us to leverage the key advantage of crowdsourcing: access to a demographically diverse participant pool that would be difficult to achieve through other recruitment methods.
@@
The study presented in this paper is part of the Emcona project, which received approval from an ethics board. Following recommendations with respect to ethical challenges in emotion analysis (Mohammad, 2022), we point out the following facts about our work: The purpose of the CAAF is to study the interplay of emotions in argument texts while taking contextual factors into account, focusing on the emotions evoked in receivers. We choose categorical emotion labels from emotion theories in accordance with previous work, and expand appraisal labels from dimensional appraisal theories to be applicable to arguments.
@@
For this study, we chose to use crowd-sourcing since the study is closer to a study in psychology than NLP. Our data collection can be seen more as an online experiment. The platform we use, Prolific, is designed for that. However, we do provide fair compensation to our study participants and report all relevant details about the online experiment. The quality of our data is ensured by multiple attention and saliency checks, which are all reported. With proper attention and saliency checks as well as a careful study design, crowdsourcing is the best option to collect lots of data from a broad variety of people with different demographics.
8.  Acknowledgements
This project has been conducted as part of the Emcona (The Interplay of Emotions and Convincingness in Arguments) project, which is funded by the German Research Foundation (DFG, project KL2869/12–1, project number 516512112). We thank Meike Bauer for her contribution to the data preparation and the fruitful discussions.
9.  Bibliographical References
[…]
@@ 9.  Bibliographical References
Appendix B Correlations
To answer our research questions, we compute correlations between all discrete emotion variables, interest, boredom, manipulation, and all appraisal variables and convincingness. The results are displayed in Table 5. The correlations are contextualized, interpreted, and analyzed in the main part of the paper.
Variable r p
Anger -0.265 0.000
Disgust -0.264 0.000
Fear 0.006 0.694
Guilt 0.105 0.000
Joy 0.435 0.000
Pride 0.458 0.000
Relief 0.511 0.000
Sadness -0.153 0.000
Shame -0.073 0.000
Surprise -0.072 0.000
Trust 0.570 0.000
Manipulation -0.084 0.000
Interest 0.693 0.000
Boredom -0.497 0.000
Suddenness -0.342 0.000
Suppression -0.326 0.000
Familiarity 0.327 0.000
Pleasantness 0.566 0.000
Unpleasantness -0.385 0.000
Consequential Importance 0.141 0.000
Positive Consequentiality 0.392 0.000
Negative Consequentiality 0.203 0.000
Consequence Manageability -0.034 0.027
Internal Check -0.103 0.000
External Check -0.355 0.000
Response Urgency -0.242 0.000
Cognitive Effort -0.061 0.000
Argument Internal Check -0.109 0.000
Argument External Check -0.497 0.000
Table 5: Pearson correlations ($r$) and corresponding $p$ values for all variables discussed in the main paper.
Appendix C Emotion Influence
@@
Figure 6: Involement question for the role-playing scenario in our annotation study.
Appendix D Topics
@@
Appendix E Role-Playing Scenario
@@
Appendix F Study Design
The following section shows the screenshots of out study as described in Section 4.3. After Figure 26 the study loops back to Figure 14 to display the second argument.